OpenStudy (anonymous):

#1: Which of the following is the mathematical representation of Boyle's law? A. P1V1 = V2/T2 B. V1/T1 = V2/T2 C. P1V1 = P2V2 D. P1/T1 = P2/T2 **my answer: C is that right?

OpenStudy (anonymous):

correct.

OpenStudy (anonymous):

#2: At room temperature, which of the following is the most compressible? A. one liter of mercury, a liquid B. one liter of oxygen, a gas C. one liter of iron, a solid D. one liter of water, a liquid **my answer: B is that right?

OpenStudy (anonymous):

ooh! yay :)

OpenStudy (chmvijay):

B. one liter of oxygen, a gas

terenzreignz (terenzreignz):

I feel I should leave the Chemistry to the Chemistry experts :D It looks like you're in good hands, @iheartfood ^.^

OpenStudy (anonymous):

#3: You open a can of soda at room temperature and hear a hiss. Which of the following factors has changed inside the container? A. pressure of gas B. temperature of gas C. type of gas D. atmospheric pressure **my answer: D is that right? :)

OpenStudy (anonymous):

@msumner i did! :) just wanted to make sure :) and @chmvijay yay :)

OpenStudy (chmvijay):

D. atmospheric pressure

OpenStudy (anonymous):

alright @terenzreignz :) haha thanks :)

OpenStudy (anonymous):

#4: Which of the following is the mathematical representation of Gay-Lussac's law? A. P1V1 = V2/T2 B. V1/T1 = V2/T2 C. P1V1 = P2V2 D. P1/T1 = P2/T2 **my answer: D is that right?

OpenStudy (anonymous):

#5: A student reduces the temperature of a 300 cm3 balloon from 60°C to 30°C. What will the new volume of the balloon be? (Remember to be careful with the units.) A. 300 cm3 B. 273 cm3 C. 150 cm3 D. 600 cm3 ***urmm not quite sure.. :/ do y'all get this?

terenzreignz (terenzreignz):

#4 is correct

OpenStudy (anonymous):

aweosme! :) and what about #5 tho? :/

terenzreignz (terenzreignz):

Well, it's Charles' Law. Volume and temperature are directly proportional. Be careful, though. Temperature must be in an absolute scale, that is, in Kelvins... Convert first.

OpenStudy (chmvijay):

one of gas law u have to apply and solve gas law at constant pressure what it is called

OpenStudy (anonymous):

yeah, but i always never understand how to do it... like how do i convert and solve? :/

OpenStudy (anonymous):

could u pls show me?

terenzreignz (terenzreignz):

Converting from Celsius to Kelvin has to to be the simplest... (compared, for instance, to Celsius to Fahrenheit, or something) To convert from Celsius to Kelvin, simply add 273.15

OpenStudy (chmvijay):

K=°C + 273

OpenStudy (anonymous):

okay... so 60+273= 333? and 30+273=303 ?

OpenStudy (anonymous):

is that right? what do i do now?

terenzreignz (terenzreignz):

Well, apply the Law... Charles' Law \[\huge \frac{V_1}{T_1}=\frac{V_2}{T_2}\]

OpenStudy (anonymous):

idk how to plug it all in :( can u pls show me?

terenzreignz (terenzreignz):

What was the volume and the temperature of the gas, initially?

OpenStudy (anonymous):

umm temp was 60 degrees and 300cm^3 was the volume... right?

terenzreignz (terenzreignz):

temperature... you will kindly express it in Kelvins :)

OpenStudy (anonymous):

okay... wait what?? the temp in kelvins? so 60 K ?

OpenStudy (anonymous):

or does that make it 333 K?

terenzreignz (terenzreignz):

More or less. Then those are your V1 and T1 respectively

OpenStudy (anonymous):

okay... so we have this so far?  or do i need to write down the units as well?

terenzreignz (terenzreignz):

No need for now. Now you also have a new temperature, that's T2

OpenStudy (anonymous):

 do i cross multiply now?